What is the RMS speed of oxygen gas at 483 K?
1
Expert's answer
2012-12-18T10:17:59-0500
& v rms = sqrt (3RT/M)
where v rms is the root mean square of the speed, M m is the molar mass of the gas in kilograms per mole, R is the molar gas constant , and T is the temperature in kelvin.
v rms = sqrt (3RT/M) = sqrt ()3*8.31 J/molK*483)/32*10^-3 kg/mol) = 613 m/s